# Active pattern formation emergent from single-species nonreciprocity

### v1

@@ Abstract
$\quad\,+\frac{B_{1}}{4\tilde{D}_{\mathrm{R}}}\left[\partial_{i}\left(\rho P_{j}\right)+\partial_{j}\left(\rho P_{i}\right)-\delta_{ij}\partial_{k}\left(\rho P_{k}\right)\right],$ (8)
with coupling coefficients $B_{i}$ ($i=1,...,5$) and the nondimensionalized rotational diffusion constant $\tilde{D}_{\mathrm{R}}$. Active Model N constitutes a minimal microscopic theory for single-species nonreciprocal interactions since its microscopic derivation (Supplementary Information) involves the minimal number of orders in the Fourier expansion [31] (namely one) required for obtaining a model that is not passive [24], and the minimal number of orders in the gradient expansion [32] (namely three) that are required for capturing the dynamics of active phase separation [23]. As shown in Methods, this model contains the previously developed Active Model B+ [28] as a limiting case, which is notable since Active Model B+ has been developed as a field theory for self-propelling particles [24].
The coupling coefficients $B_{i}$ can be explicitly calculated as moments of the microscopic nonreciprocal forces and torques. They vary with the vision-cone opening angle $\vartheta$ and the force and torque parameters. Detailed expressions are given in equations (18)–(22) of Methods for the choice
@@
which describes a force with a short-ranged repulsive contribution of strength $F_{0}$ and a long-ranged attractive contribution of strength $F_{1}$, and a torque of strength $b$. The length scales $a$ and $c$ determine the range on which the forces and torques act. In the following we use this choice of soft interaction, and set weaker strength of attractive force (e.g., $F_{1}/F_{0}=10^{-3}$) and torque ($b/(F_{0}a)=0.3$) as compared to the repulsive force, but with longer interaction range ($\alpha=c/a=2$). Note that the nonreciprocal active field theory given by equations (6) and (7) is generic and the forms of model terms do not depend on the specific choices of force and torque functions. The combination of a short-ranged repulsive and a long-ranged attractive force holds for many interactions usually considered.
Bifurcations and phase diagram of active patterns
@@
where $B_{5}$ is given by equation (22) (Methods), when
@@
A primary bifurcation analysis can be conducted for the fully disordered state (Supplementary Information). For the soft interaction forms (9) chosen here, it yields a flocking transition (without phase separation) when equation (11) is satisfied at large enough average particle density $\bar{\rho}$ or opening angle $\vartheta$, with lower transition threshold for larger torque strength $b$ and range $c$, consistent with the torque aligning effect. The resulting homogeneous flocking phase is characterized by uniform distribution of particle density ($\rho=\bar{\rho}$) with constant polarization alignment ($\mathbf{P}=\bar{\mathbf{P}}\neq\mathbf{0}$), i.e., a uniform aligned phase where the orientational symmetry is spontaneously broken. More interesting transitions occur for the subsequent secondary bifurcation from this orientationally symmetry-breaking state, with the linearized dynamical equation for Fourier components of perturbations $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ given by
@@
where details of the dynamical matrix $\mathcal{L}$ are presented in Supplementary Information. When $\vartheta\neq\pi$ (with a limited vision perception $2\vartheta<2\pi$), $\mathcal{L}^{\dagger}\neq\mathcal{L}$, i.e., $\mathcal{L}$ is non-Hermitian, leading to nonreciprocal, parity-time (PT) symmetry breaking phase transitions [14, 16]. Note that here the system non-Hermiticity originates from vision-cone nonreciprocal interactions within single species, fundamentally different from the previous cases featured by asymmetric coupling between different populations (e.g., interspecies cross diffusivities) [15, 14, 17, 16] or different harmonic modes [16].
@@
Fig. 2: Active patterns and phase diagram. a-e Simulation snapshots for various active patterns that emerge at average particle density $\bar{\rho}=1.5$, including active stripes at $\vartheta=30^{\circ}$ (a and Supplementary Video 1), active branches at $\vartheta=90^{\circ}$ (b and Supplementary Video 2), active yarn at $\vartheta=140^{\circ}$ (c and Supplementary Video 3) and $160^{\circ}$ (d and Supplementary Video 4), and a homogeneous flocking phase with vortex defects of polarization field $\mathbf{P}$ at the reciprocal limit $\vartheta=180^{\circ}$ (e and Supplementary Video 5). In a-d with patterns induced by nonreciprocity, the spatial profiles of particle density $\rho$ are presented, and the local distribution of vector field $\mathbf{P}$ are indicated as arrows in the insets which are enlarged portions of the boxed regions. The patterns self-travel opposite to the average direction of $\mathbf{P}$ (Supplementary Videos 1-5). In e the spatial profile of polarization magnitude $|\mathbf{P}|$ is shown. f The phase diagram of $\bar{\rho}$ vs $\vartheta$, with solid curves evaluated from the analytical results of bifurcation analysis and the symbols identified via outcomes of numerical simulations (those giving disordered or homogeneous phases are not shown).
@@
We have calculated the corresponding $\bar{\rho}$ vs $\vartheta$ phase diagram, as shown in Fig. 2, by evaluating the analytical results of bifurcation analyses (Supplementary Information), with phase boundaries well agreeing with numerical simulations of the full model equations (6)–(8). A noteworthy phenomenon is the occurrence of phase transitions with the increase of $\vartheta$ or $\bar{\rho}$, i.e., from (i) the disordered state of both particle and polarization densities, to (ii) a traveling striped state of low- vs high-density microphase separation and local flocking alignment (Fig. 2a and Supplementary Video 1), then to (iii) homogeneous flocking with again disordered particle density, and back to (iv) a phase separated, flocking state with emergent dynamical chiral patterns (i.e., patterns for which spatial mirror symmetry is broken), including active branches (Fig. 2b and Supplementary Video 2) and active yarn pattern which shows as a phase consisting of traveling bands containing microphase-separated fine texture that are embedded in a homogeneous medium (see Fig. 2c,d and Supplementary Videos 3, 4), before (v) a reentrance to the homogeneous flocking phase at $\vartheta=\pi$. At the limit of $\vartheta=\pi$ with complete perception, the system becomes reciprocal and the dynamics of $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ are decoupled, resulting in the stability of uniform flocking with homogeneous $\rho$ and the usual vector field dynamics for polarization $\mathbf{P}$ with motion and annihilation of vortex defects (Fig. 2e and Supplementary Video 5). The phase separation effects observed here are thus induced by vision-cone-type force nonreciprocity.
@@
Although no explicit self-driving factor has been incorporated in the microscopic model, all phase-separated patterns self-travel unidirectionally, which leads to the breaking of time-reversal symmetry. The direction of pattern traveling is opposite to that of the average polarization field $\langle\mathbf{P}\rangle$, with average migration velocity $\bar{\mathbf{v}}\propto-\langle\mathbf{P}\rangle$. This is consistent with the microscopic picture: Due to nonreciprocity, each particle can only interact with the neighboring particles within its own vision cone centered around its orientation (Fig. 1). Since in the setup here the short-range repulsion is much stronger than the long-range attraction, the particle feels the net effect of repulsion from the particles in front of it (within the vision cone) but not from those behind, such that every particle ends up being pushed towards the direction opposite to its orientation. Given that $\mathbf{P}$ measures the particle orientation, collective particle migration occurs if equation (11) is satisfied with strong enough torque alignment. As particles are still being attracted to next neighbors, such a competing effect leads to the occurrence of phase separation at large enough particle density or opening angle. This phenomenon of nonreciprocity-induced phase separation with local orientational alignment then incorporates both types of collective behavior of active particles [33, 34], assembly via phase separation, and aligning collective self-migration, with mechanisms distinct from the previously known motility-induced [35] or nonreciprocal-torque-based [33] phase separation.

### v2

@@ II.3 Nonreciprocal field theory: Active Model N
$\quad\,+\frac{B_{1}}{4\tilde{D}_{\mathrm{R}}}\left[\partial_{i}\left(\rho P_{j}\right)+\partial_{j}\left(\rho P_{i}\right)-\delta_{ij}\partial_{k}\left(\rho P_{k}\right)\right]$ (13)
with coupling coefficients $B_{i}$ (i=1,,5i=1,...,5) and the nondimensionalized rotational diffusion constant $\tilde{D}_{\mathrm{R}}=a^{2}D_{\mathrm{R}}/D_{\mathrm{T}}$. Active Model N developed here constitutes a minimal microscopic theory for single-species nonreciprocal interactions since its microscopic derivation (Appendix A) involves the minimal number of orders in the Fourier expansion [46] (namely one) required for obtaining a model that is not passive [27], and the minimal number of orders in the gradient expansion [47] (namely three) that are required for capturing the dynamics of active phase separation [26]. As shown in Appendix B, this model contains the previously developed Active Model B+ [32] as a limiting case, which is notable since Active Model B+ has been developed as a field theory for self-propelling particles [27]. This is in line with similar observations made recently for reaction-diffusion systems [48] and shows that theories like Active Model B+ are significantly more general than what specific microscopic derivations might suggest.
The coupling coefficients $B_{i}$ can be explicitly calculated as moments of the microscopic nonreciprocal forces and torques. They vary with the vision-cone opening angle $\vartheta$ and the force and torque parameters, as shown in Appendix A. In this study we choose the specific form of force and torque functions as
@@
which describes a force with a short-ranged repulsive contribution of strength $F_{0}$ and a long-ranged attractive contribution of strength $F_{1}$, and a torque of strength $b$. The length scales $a$ and $c$ determine the range on which the forces and torques act. After rescaling over the length scale $a$, the nondimensionalized coefficients $B_{i}$ in Eqs. (11) and (12) are expressed as
$B_{1}$ $=\pi\beta a\left(F_{0}-2\alpha^{2}F_{1}\right)\sin\vartheta,$
$B_{2}$ $=\frac{\pi}{2}\beta a\left(\sqrt{\pi}F_{0}-8\alpha^{3}F_{1}\right)\vartheta,$
$B_{3}$ $=\frac{\pi}{8}\beta a\left(F_{0}-12\alpha^{4}F_{1}\right)\sin\vartheta,$
$B_{4}$ $=\frac{3}{32}\pi\beta a\left(\sqrt{\pi}F_{0}-64\alpha^{5}F_{1}\right)\vartheta,$
$B_{5}$ $=-2\pi\beta b\left(\frac{c}{a}\right)^{2}\vartheta.$ (15)
In the following, we use this choice of soft interaction and set a weaker attractive force (e.g., $F_{1}/F_{0}=10^{-3}$) and torque ($b/(F_{0}a)=0.3$) relative to the repulsive force, but with a longer interaction range ($\alpha=c/a=2$). The corresponding expressions of rescaled model parameters are summarized in Appendix D. Note that the nonreciprocal active field theory given by Eqs. (11) and (12) is generic and the forms of model terms do not depend on the specific choices of force and torque functions (see Appendix A). The combination of a short-ranged repulsive and a long-ranged attractive force holds for many interactions usually considered.
It is also important to note that in the model constructed here, only nonreciprocal interactions are incorporated (with the use of Eqs. (4) and (5)), such that all the observed effects are solely a consequence of single-species nonreciprocity. Of course, actual physical or biological living systems are governed by both nonreciprocal and reciprocal interactions (such as reciprocal steric repulsion that comes into play at very short distances) as well as self-propulsion. Thus, the effects and phenomena observed in our model will not directly map onto those observed in, e.g., real living systems. Nevertheless, they provide fundamental insights into novel physical phenomena that, in a more complex form, are relevant for living systems; also importantly, the resulting predictions can be examined in experimentally controllable artificial systems such as specifically-designed colloidal particles [3]. Moreover, it is quite common in biological systems that activity takes the form not of self-propulsion of individual agents but of active relative motion, and the study of such phenomena can benefit substantially from microscopically derived active field theories, as demonstrated recently for microtubule-motor mixtures interacting via sliding and alignment [49].
III Bifurcations and phase diagram of active patterns
@@
where $B_{5}$ is given by Eq. 38, when
@@
A primary bifurcation analysis can be conducted for the fully disordered state (Appendix C). For the soft interaction forms of Eq. 14 chosen here, it yields a flocking transition (without phase separation) when Eq. 17 is satisfied at large enough average particle density $\bar{\rho}$ or opening angle $\vartheta$, with lower transition threshold for larger torque strength $b$ and range $c$, consistent with the torque aligning effect. The resulting homogeneous flocking phase is characterized by uniform distribution of particle density ($\rho=\bar{\rho}$) with constant polarization alignment ($\mathbf{P}=\bar{\mathbf{P}}\neq\mathbf{0}$), i.e., a uniform aligned phase where the orientational symmetry is spontaneously broken. More interesting transitions occur for the subsequent secondary bifurcation from this orientationally symmetry-breaking state, with the linearized dynamical equation for Fourier components of perturbations $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ given by
@@
where details of the dynamical matrix $\mathcal{L}$ are presented in Eq. 135 of Appendix C. When $\vartheta\neq\pi$ (with a limited vision perception $2\vartheta<2\pi$), $\mathcal{L}^{\dagger}\neq\mathcal{L}$, i.e., $\mathcal{L}$ is non-Hermitian, leading to nonreciprocal, PT symmetry breaking phase transitions [18, 20]. Note that here the system non-Hermiticity originates from vision-cone nonreciprocal interactions within single species, fundamentally different from the previous cases featured by asymmetric coupling between different populations (e.g., interspecies cross diffusivities) [19, 18, 21, 20] or different harmonic modes [20].
@@
Figure 2: (a)–(e) Simulation snapshots for various active patterns that emerge at average particle density $\bar{\rho}=1.5$, including active stripes at $\vartheta=30^{\circ}$ ((a) and Supplemental Video 1), active branches at $\vartheta=90^{\circ}$ ((b) and Supplemental Video 2), active yarn at $\vartheta=140^{\circ}$ ((c) and Supplemental Video 3) and $160^{\circ}$ ((d) and Supplemental Video 4), and a homogeneous flocking phase with vortex defects of polarization field $\mathbf{P}$ at the reciprocal limit $\vartheta=180^{\circ}$ ((e) and Supplemental Video 5). In (a)–(d) with patterns induced by nonreciprocity, the spatial profiles of particle density $\rho$ are presented, and the local distribution of vector field $\mathbf{P}$ are indicated as arrows in the insets which are enlarged portions of the boxed regions. The patterns self-travel opposite to the average direction of $\mathbf{P}$ (Supplemental Videos 1-5). In (e) the spatial profile of polarization magnitude $|\mathbf{P}|$ is shown. (f) The phase diagram of $\bar{\rho}$ vs $\vartheta$, with solid curves evaluated from the analytical results of bifurcation analysis and the symbols identified via outcomes of numerical simulations (those giving disordered or homogeneous phases are not shown).
@@
We have calculated the corresponding $\bar{\rho}$ vs $\vartheta$ phase diagram, as shown in Fig. 2, by evaluating the analytical results of bifurcation analyses (Appendix C), with phase boundaries well agreeing with numerical simulations of the full model equations (11)–(13). A noteworthy phenomenon is the occurrence of phase transitions with the increase of $\vartheta$ or $\bar{\rho}$, i.e., from (i) the disordered state of both particle and polarization densities, to (ii) a traveling striped state of low- vs high-density microphase separation and local flocking alignment (Fig. 2(a) and Supplemental Video 1), then to (iii) homogeneous flocking with again disordered particle density, and back to (iv) a phase separated, flocking state with emergent dynamical chiral patterns, i.e., patterns for which spatial mirror symmetry is broken, including active branches (Fig. 2(b) and Supplemental Video 2) and active yarn pattern which shows as a phase consisting of traveling bands containing microphase-separated fine texture that are embedded in a homogeneous medium (see Figs. 2(c), 2(d) and Supplemental Videos 3, 4), before (v) a reentrance to the homogeneous flocking phase at $\vartheta=\pi$. In the limit of $\vartheta=\pi$ with complete perception, the system becomes reciprocal and the dynamics of $\widehat{\rho}_{\mathbf{q}}$ and $\widehat{\mathbf{P}}_{\mathbf{q}}$ are decoupled, resulting in the stability of uniform flocking with homogeneous $\rho$ and the usual vector field dynamics for polarization $\mathbf{P}$ with motion and annihilation of vortex and antivortex defects (Fig. 2 (e) and Supplemental Video 5). The phase separation effects observed here are thus induced by vision-cone-type force nonreciprocity.
The active branches and active yarn patterns are distinguished through two main characteristics, (i) microphase separation with the formation of spatially dispersed fibers for active branches vs the simultaneous occurrence of micro and bulk phase separations for the dense banding of active yarn, and (ii) the appearance of different fine textures that are dominated by bifurcated microstructures in active branches but multi-branching in active yarn, as will be further detailed in the next section and Fig. 4. Both types of patterns vary persistently with time and are the dynamical results of nonlinear evolution far from equilibrium; thus the phase boundary between them (see Fig. 2(f)) cannot be determined by linear stability or bifurcation analysis but was estimated through simulations. With the increase of opening angle $\vartheta$, around the boundary of transition the patterns start to develop large gaps of homogeneous medium separating clusters/bands of fibers (instead of spreading throughout the whole system), with the mixing of both bifurcated and multi-branched fiber textures.
Figure 3: (a) Visualization of a single particle with orientation $\hat{\mathbf{u}}$ and vision-cone opening angle $2\vartheta$. Via interacting with other nearby particles within its field of vision, it experiences strong short-range repulsive (red) and weak long-range attractive (green) forces, as well as aligning torques (blue). The individual forces and torques (light arrows) result in a net force and a net torque (bold arrows). (b) On a microscopic level, spontaneous motion originates from the net repulsive forces in the bulk. Each particle experiences a net repulsion by the particles in front of it (within its vision cone) but no force from those behind, leading to collective motion in the direction of the bold violet arrow on the right. (c) Visualization of the microscopic mechanism for branch merging. If two branches cross, the particles belonging to different branches start seeing each other and, due to the aligning torques, acquire the same orientation. As a result, they travel within one merged branch.
@@
Although no explicit self-driving factor has been incorporated in the model, all phase-separated patterns self-travel unidirectionally, which leads to the breaking of time-reversal symmetry. The direction of pattern traveling is opposite to that of the average polarization field $\langle\mathbf{P}\rangle$, with average migration velocity $\bar{\mathbf{v}}\propto-\langle\mathbf{P}\rangle$. This is consistent with the microscopic picture: Due to nonreciprocity, each particle can only interact with the neighboring particles within its own vision cone centered around its orientation (Fig. 1). Since in the setup here the short-range repulsion is much stronger than the long-range attraction, the particle feels the net effect of repulsion from the particles in front of it (within the vision cone) but not from those behind, such that every particle ends up being pushed towards the direction opposite to its orientation, as illustrated in Figs. 3(a) and 3(b). Given that $\mathbf{P}$ measures the particle orientation, collective particle migration occurs if Eq. 17 is satisfied with strong enough torque alignment. As particles are still being attracted to next neighbors, such a competition between short- and long-range interactions leads to phase separation at large enough particle density or opening angle. This phenomenon of nonreciprocity-induced phase separation with local orientational alignment then incorporates both types of collective behavior of active particles [50, 51], assembly via phase separation, and aligning collective self-migration, with mechanisms distinct from the previously known motility-induced [52] or nonreciprocal-torque-based [50] phase separation.